Known: The mass and the acceleration of gravity are givern Find: Determine the weight of the object in: (a) English Engineering units (b) British Gravitational units (c) SI units Schematic and Given Data: 9.81 m/s2 10 kg
Answer #1

1) m = 10Kgs

g= 9.8m/sec^2

c) SI Unit

W= m*g = 10*9.8 = 98N

a) EnglishEngineering unit

m= 10*2.204lb = 22.04 lb

g = 32.174ft/sec^2

W = mg = 22.04*32.174 = 709.11 lbf

b)British Gravitation unit

W = 709.11lbf

2)mg = 1N

del x = 5mm

k = mg/del x = 1/5 N/mm

Now del x1 = 30mm

k = Mg/del x1

1/5 = M*9.8/30

M = 30*9.8/5 = 58.8 Kg